OpenStudy (anonymous):

An object A has half the mass and twice the velocity of object B. How are the kinetic energies of A and B related? A. The kinetic energy of A is the same as that of B. B. The kinetic energy of B is twice that of A. C. The kinetic energy of A is twice that of B. D. The kinetic energy of B is 4 times that of A.

OpenStudy (anonymous):

B = (1/2) M V^2 A = (1/2) (1/2 * M) (2V^2) plug in an example like M = 2 V = 4 B = (1/2) (2) (16) = 16J A = (1/2) (1/2 * 2) (2*4^2) = (1/2) (1) (32) = 16J so the answer is A right @JoannaBlackwelder

OpenStudy (joannablackwelder):

Close. The 2V needs to be squared, not just the V.

OpenStudy (anonymous):

oh so it's A = (1/2) (1) (64) = 32J so it's C?

OpenStudy (joannablackwelder):

Looks right to me :)

OpenStudy (anonymous):

woooo thank you!